OpenStudy (photon336):

A magazine ranks companies based on their overall percentage revenue growth over the previous three years. Last year, company A was ranked 5000 and had revenues that were 134% of what they were three years prior. By what percentage will you need to increase revenues per year to meet the mark set by company A (34% growth rate in three years) in order to give your company a good chance of making the list three years from now. (product of n numbers)^(1/n)-1

OpenStudy (photon336):

I placed the formula I need to apply below, I'm a bit confused as to what the formula "means" as it applies to this problem

OpenStudy (michele_laino):

I'm sorry I'm not good with financial mathematics

OpenStudy (photon336):

okay :)

OpenStudy (photon336):

@mayankdevnani

OpenStudy (photon336):

@Kainui @Mehek14

OpenStudy (phi):

where is the formula ?

OpenStudy (photon336):

(Product of n numbers)^(1/n) - 1 this was what I was given

OpenStudy (phi):

also, is there a typo in this? I can't make sense of company A had revenues that were 134% of what they were three years prior. and the mark set by company A (34% growth rate in three years). is that supposed to be 134% growth rate ??

OpenStudy (phi):

I would think about this way. we assume we grow at a rate, for example, 10% or as a decimal 0.1 that means if we start with x and grow 10% we do: x+ 10% of x or as a decimal: x+0.1x if we "factor out" the x, we can write it as x(1+0.1) x is the amount we started with and after 10% grow we end up with x*1.1 if we don't know the rate, then let's give it a name "r" x(1+r) after the 2nd year: x(1+r)(1+r) and the 3rd year: x(1+r)^3 and that equal x plus 134% of x or x+1.34x= 2.34x so we have \[ x(1+r)^3 = 2.34x\] divide by x \[ (1+r)^3 = 2.34\] now take the cube root of both sides 1+r= (2.34)^(⅓) finally r= (2.34)^(⅓)-1 change the decimal to a percent for the final answer